### v3

@@ Abstract
Automatic markerless estimation of infant posture and motion from ordinary videos carries great potential for movement studies ”in the wild”, facilitating understanding of motor development and massively increasing the chances of early diagnosis of disorders. There is rapid development of human pose estimation methods in computer vision thanks to advances in deep learning and machine learning. However, these methods are trained on datasets that feature adults in different contexts. This work tests and compares seven popular methods (AlphaPose, DeepLabCut/DeeperCut, Detectron2, HRNet, MediaPipe/BlazePose, OpenPose, and ViTPose) on videos of infants in supine position and in more complex settings. Surprisingly, all methods except DeepLabCut and MediaPipe have competitive performance without additional finetuning, with ViTPose performing best. Next to standard performance metrics (average precision and recall), we introduce errors expressed in the neck-mid-hip (torso length) ratio and additionally study missed and redundant detections, and the reliability of the internal confidence ratings of the different methods, which are relevant for downstream tasks. Among the networks with competitive performance, only AlphaPose could run close to real time (27 fps) on our machine. We provide documented Docker containers or instructions for all the methods we used, our analysis scripts, and the processed data at https://hub.docker.com/u/humanoidsctu and https://osf.io/x465b/.
1 Introduction
@@
Methods for automatic human keypoint extraction (for example, eyes, wrists, hips, feet, etc.), also called human pose estimation, from images and videos are rapidly evolving, with performances increasing every year (see [34, 35, 36] for recent surveys). These provide key enabling technology to make automatic motion extraction and analysis from “in the wild” recordings possible. These methods were developed primarily to estimate posture and movement of adult bodies (see [37] for a recent comparison of OpenPose and marker-based motion capture). So far, these methods have been applied to infant videos “as is”. However, the morphology of the infant body is different from the body proportions of the adult, especially in early infancy [38, 39]. Additionally, small infants are typically in supine position (on their back) and move differently than standing adults, which constitutes a dataset which is out of the distribution the adult pose estimation models were trained on. Although automated pose estimation seems to be generally good (see Fig. 1 a-c), in a previous study [11], some limitations of current pose estimation methods in infants were highlighted. For example, OpenPose [28] was found to struggle with some camera angles, body postures, and keypoints. In particular, complex leg positions, such as when legs are crossed in supine position, are not often estimated with high accuracy. This can be seen in Fig. 1 (d): OpenPose misses the left ankle keypoint and places the left knee keypoint on the thigh, very close to the left hip. In the current study, comparing different pose estimation methods, we found similar examples. Fig. 1 (e) shows that HRNet Bottom-Up fails on a complex leg position and fails to place the left shoulder and elbow keypoints, placing them on the right; while (b) shows that HRNet Top-Down hallucinates a second person in the image (see the additional bounding box at the infant’s right forearm), and (f) shows ViTPose not detecting the infant while sitting on the lap of a parent (the image was cropped to not show the parent’s face in the figure).
@@
Figure 1: Examples of successful and erroneous keypoint estimations obtained from the different human pose estimation methods: a) HRnet Bottom-Up (”supine” dataset, infant AA, 17 weeks) successful, b) HRnet Top-Down (”supine” dataset, TH, 10 weeks) with a successful estimation but a second hallucinated person, c) ViTPose (SyRIP image 293) successful, d) OpenPose missing the left ankle keypoint (synthetic infant 1); e) HRNet Bottom-Up failing to estimate a complex leg position (synthetic infant 10); f) ViTPose detecting the parent but not the infant (”Lap” dataset).
@@
Thus, our first goal and contribution is to assess the performance of 2D human pose estimation methods when applied specifically to infants in supine position. In order to provide a more general outlook, we additionally present results on images of infants in a wider variety of postures, including standing or turning away from the camera, as well as with infants on the lap, or sitting or standing in front, of their parent. Only few studies have partially addressed this problem. However, if present, “in the wild” recordings did not constitute the majority of the datasets. Needham et al. [19] used a multi-camera setup in the lab and compared OpenPose, AlphaPose and DeepLabCut. Groos et al. [40] created a large proprietary dataset of human-annotated infant images from both controlled and standardized environments and “in the wild” recordings, and used them to retrain several models, including OpenPose. Their best model achieves on average 81.11% of its estimates being located within an error of 10% of the length of the infant’s face, while the “vanilla” OpenPose had 49.66% of its estimates within this margin of error. Sermpon et al. [41] have found weak correlations between pose estimation features derived from MediaPipe’s estimates and manually annotated movements towards the midline. The preprint of Jahn et al. [42] compared OpenPose, MediaPipe, HRNet, AggPose and VitPose, as well as fine-tuned versions of ViTPose and HRNet (AGMA-HRNet48) with infant recordings in a laboratory context made to specifically target GMA settings (infants in supine position, seen from above and at a diagonal angle). Their results show that ViTPose-huge has the second best results (higher than AGMA-HRNet48, the HRNet model fine-tuned with infant images from a similar GMA context), locating 84.6% of its estimates within 2.5cm of the ground truth, while their re-trained ViTPose locates 93.89% of estimates within this distance.
@@
In this article, we compare the performance of seven methods: Alphapose [31], DeepLabCut / DeeperCut (v2.2.1.1) [43], Detectron2 (v0.6) [44], MediaPipe / BlazePose (v0.10.14) [45, 46], HRNet (MMPose v0.28.0) [47, 48] (Bottom-Up (HRnet BU) and Top-Down (HRnet TD)), OpenPose [28] and ViTPose (MMPose v1.1.0) [49].
@@
As a second contribution of this work, we put together and evaluated a number of measures to assess the performance of the pose estimation methods. Besides the ones typically used by the machine learning community (Average Precision and Recall), we consider alternative measures that express the errors scaled to the infant’s body dimensions (Neck-MidHip error), and we additionally analyze the percentage of missed and redundant detections and the reliability of the confidence estimates that the pose estimation methods output. We also report the processing time of the methods. All of these together provide a more complete picture and allow the user to make an optimal choice of the method.
@@
The third contribution of this work constitutes in making all the pose estimation methods and tools for the results analysis available to the community. Although the pose estimation algorithms themselves are publicly accessible in their corresponding repositories, we make available the versions we used including the complete environment in public Docker containers (at https://hub.docker.com/u/humanoidsctu) with instructions to simplify their use. In addition, we also share the evaluation scripts and the detailed results (at https://osf.io/x465b/).
@@ Methods
To ensure that comparisons are as fair as possible, we used versions of the pose estimation methods that were trained with the same dataset: COCO [50], although some methods only have available weights that used other datasets in their training, instead of or in addition to COCO (see each method’s respective subsection below or the summary Tab. 1). The methods provide 17 estimated keypoints, with the exception of DeepLabCut, MediaPipe and OpenPose that provide 14, 33 and 18 keypoints respectively. We will call a single set of estimated keypoints a detection. The details of the parameters with which the methods were run are provided in the supplemental materials. All methods except MediaPipe provide a form of confidence values as their own internal evaluation of the estimated result, with higher values indicating higher confidence in the quality of the estimates. This value can be provided for the whole set of keypoints or for individual keypoints (or both). Following the trend in the documentation and methods’ outputs, we will use score when referring to the whole detection value and confidence when referring to the individual keypoints values. The methods follow either one of the two approaches: Top-Down, which first uses a detector to delimit areas of interest in the image that encompass a person to detect and then estimates a set of keypoints within this space; and Bottom-Up, which finds keypoints or sets of keypoints separately on the whole image before joining them into full keypoint sets and separating them into several individuals if necessary.
@@
  • Real infants supine dataset. 720 annotated images (90 per video) from 8 videos (59200 images total) of two infants, followed longitudinally between 2 and 6 months of age. The annotations were made using the DeepLabCut labeling tool. 480 images (60 per video) were chosen by taking the first ten images every 100 images, from 0 to 9, then 100 to 109, until 509. The remaining 240 images (30 per video) were selected through DeepLabCut’s integrated selection tool using k-means clustering to pick the least similar images in each video. Fourteen keypoints were annotated, corresponding to the COCO keypoints minus the ears and the nose. Our dataset is entirely comprised of infants recorded at home by their parents, each video being different from the others in terms of camera angles, lighting conditions, as well as background and clothes colors and type.
  • Synthetic infants – MINI-RGBD. Based on recordings of real infants, Hesse and colleagues have trained a model capturing infant shape and posture distributions which they subsequently used to produce and render synthetic sequences of artificial infants under 7 months in supine position—the MINI-RGBD dataset [54]. There are 12 synthetic infants with 1000 images each. Ground truth for 25 keypoints is available. Of these 25 keypoints, 13 are in common with COCO’s 17 keypoints and were used to evaluate the methods. Each of the twelve synthetic sequences is unique, including backgrounds with large variations between sequences (see Fig. 1 for two examples, or see the original paper [54]).
@@
  • Real infants on the lap dataset. 359 annotated images from 6 videos of 6 different infants aged from 17 weeks to 18 months. Two of these videos (infants) were recorded at home by the parents, with the infant sitting on the lap of a parent, themselves sitting on a couch. The remaining four videos (infants) were recorded in a laboratory context, with the camera placed in front of the participants, who were sitting (or standing) on a mat placed on the ground. The parent, and in particular their face, was not always fully visible in the images. The number of 359 images was reached after filtering down the images that were originally annotated but where the infant was missing (moved out of the field of the camera), or because an experimenter was in front of the camera and occluded most of the infant.
@@
For real infants, it is difficult to estimate the difficulty of each sequence or image, though it is clear that our “lap” and the SyRIP datasets would be more difficult to estimate than our ”supine” dataset. The SyRIP dataset contains a wider variety of infant postures and camera angles, including many that partially hide the infants (e.g., from the side), making the estimation more difficult. As for the ”lap” dataset, since the methods were all originally trained on datasets of adult images, they are biased toward detecting adults, and the presence of the adult overlapping right behind the infant is a strong source of disturbance. For synthetic infants, the MINI-RGBD dataset has a definition of the sequence difficulty: easy (IDs 1 to 4), medium (5 to 9) and difficult (10 to 12), with this difficulty judged primarily with regard to the posture of the infant. Although the main manuscript focuses on the average results, it might be interesting to look at the results at the level of individual videos or their estimated difficulty. Such results are available in the Supplementary Materials.
@@
Our supplementary materials include the results for an extended version of our ”supine” dataset: 720 supplementary images from 8 additional videos; as well as a continuous sequence of 900 images (30 seconds) of a single infant at 17 weeks old. A summary table of each dataset’s characteristics is available in Sup. Tab. 1.
@@
The metrics used for comparisons are described below. Because we observed differences in the results when the methods were given input in the form of videos or images, we present the results for both input types independently, except for DeepLabCut and ViTPose (within the MMPose environment) that only accepted video input.
@@
Out of all these metrics, the ones that can be used even without ground truth, and for which we have the results on the full videos on our ”supine” dataset are: percentages of missing data and redundant detections, and the processing speed.
@@
To evaluate the performance of pose estimation, the gold standard metrics in the literature are Average Precision and Average Recall [56], ranging between 0 and 100, higher values indicating better performance. We follow the benchmark evaluation set by the COCO challenge to calculate AP and AR, averaging the values computed over ten thresholds from 0.5 to 0.95, with steps of 0.05. Roughly, if the OKS for a detection is above the threshold and there is any ground truth that has not been matched with a detection yet on a given image, then the detection counts as True Positive. Other cases count as False Positives. If there is no detection but there is a ground truth, then it counts as a False Negative. Precision and Recall are computed for each threshold, then averaged. Note that despite the name Average Precision, it is actually computed as the area under the Precision-Recall Curve[57, 56]. In our case, the computation is simplified, as all our annotated images contain one single infant (one ground truth). For our ”Lap” dataset, the ground truth annotations of the parents were ignored, so that each method was only evaluated against the infant annotations.
@@
As an alternative to OKS, we additionally use a reference distance, the Neck-MidHip length, corresponding roughly to the length of the infants’ torso. With this reference distance, we normalize the errors between recordings to compensate for different camera settings and infants of different ages and with different body sizes. Others and ourselves have previously used a similar strategy in a similar context [29, 11, 42]. Unlike OKS, such errors are easier to grasp as they directly relate to a measurable part of the body of the infant. Lower is better, and the minimum is 0.
The equations are shown in Eq. 2. Given $j$ a recording with $I_{j}$ annotated images for which Neck and MidHip keypoints ground truth are available or can be approximated. Given the images from the recording $i\in[0,I_{j}]$, then the Neck-MidHip length for this image $nmh\_len_{ji}$ is defined in pixels as the Euclidean distance between the Neck and the MidHip ground-truth coordinates in this image. For our ”supine” and the MINI-RGBD dataset, we take the median of all $nmh\_len_{ji}$ as the NeckMidHip length to use to normalize the recording: $nmh\_len_{j}$. Given a keypoint $k$ that is visible in the image and annotated, and detected by the pose estimation method, then its Neck-MidHip error $nmh\_error_{kij}$ is its error (Euclidean distance) relative to its ground truth, $d_{kij}$, divided by the median Neck-MidHip length for the recording, $nmh\_len_{j}$. For our ”lap” and the SyRIP dataset, which contain images from different sources mixed together, we normalize using each images’ neck-MidHip distance $nmh\_len_{ji}$ instead of a median from a whole recording. This has the downside that only images where this length can be estimated are used to calculate the error. Then, we obtain per-keypoint-location mean Neck-MidHip errors, and an overall mean full-body error, from all images and recordings.
@@
For the synthetic infants in the MINI-RGBD dataset, the Neck and MidHip positions are provided directly in the ground truth. In the real infants datasets, the ground truths do not provide Neck or MidHip keypoints. Instead, the centre of the right and left Shoulder and Hip keypoints are used, turning into a “MidShoulder-MidHip” length. We will simply write Neck-MidHip for simplification in the rest of the paper. The Neck-MidHip length is computed using the median of the Neck-MidHip lengths for a whole video when possible. This enables the use of all annotated images and keypoints even if one of the keypoints needed to compute the length is missing from the ground truth in a particular image (occlusions). It is also less subject to noise in the annotations, with the drawback that a single length is used for a whole video and does not fit exactly to each individual image. Note that using the median can only be done when there are no changes in camera settings, in particular zoom or camera position, which can happen, for example, when recordings are performed at home by a parent holding a smartphone with their hand; otherwise, it requires to separate the recording into subsets according to the different settings, or use per-image normalization instead, which is what we opted to do for our ”lap” and the SyRIP datasets.
@@
The total percentage of missing data computation is described in Eq. 3. We define the maximum amount of data $maxdata_{m}$ that can be estimated by a method $m$ as the number of images to estimate $N$ times the number of keypoints provided by the method $m_{kp}$. Missing detections $mdet$ are converted to this unit by multiplying with $m_{kp}$ as well, since any missing detection is equivalent to missing all keypoints in this one image. This number is added to the count of individual missing keypoints $mkp$, and divided by $maxdata_{m}$ to obtain the total percentage of missing data for this method: $mdata_{m}$.
@@
Some methods can erroneously detect more people than there are in the image, as shown in Fig. 1 (b), where an additional ”person” is detected. These redundant detections might indicate that the method hallucinates other people on the image where there are none, or provides several estimations for the same person with different positions. Such behavior is undesirable, in particular for applications with more complex environments that, for example, involve interaction with an experimenter or the parents, such as in our ”lap” dataset, as it complexifies tracking each separate individual when ground truth is not available. The percentage of redundant detections is computed as the ratio between i) the difference in the number of detections between the expected and provided detection numbers (number of estimated detections minus the number of expected detections) divided by ii) the number of images the method provided detections for (total number of images minus the number of missed detections). For example, if there are 150 detections for 100 input images with 100 expected detections (1 infant per image), then 50 detections would be considered redundant, resulting in a percentage of 50% of redundant detections, translating that, on average, the method provides one additional detection every two images where it detected something. A percentage of 100% would mean that, on average, the method always provides two detections for each image, when one is expected. DeepLabCut, MediaPipe, and OpenPose did not have redundant detections (MediaPipe and OpenPose, have a built-in parameter to limit the detections to 1 person per image).
In the case of the ”Lap” dataset, where an adult is always present behind the infant, there are also cases where only the adult and not the infant is detected (Fig. 1 (f)). We empirically observed that some pose estimation methods have a tendency to often detect only the adult and not the infant in such context. This would show as a negative ratio of redundant detections. Taking the example from above, if a method had provided 150 detections for 100 images where 200 detections were expected due to the presence of both an infant and an adult, then we would get -50% “redundant” detections: on average, that method provided the two expected detections for half of the images, and only provided one detection for the other half. For this particular dataset, this is not a foolproof measure, as it is possible that a method detected the adult and then provided a second, redundant, detection of the adult or hallucinated another person somewhere else in the image instead of detecting the infant.
@@
In a real application scenario, without available ground truth, confidence or score values can be useful to provide cues about detections or keypoints that should not be trusted. For example, confidences are used as coefficients in the optimization process of smplify-x [58] when finding the parameters of the 3D full-body pose estimation from the 2D pose estimation. These values should be positively correlated with the OKS (the higher the OKS, the higher the score). When available, we use the score provided by the methods for each image. When not available, we used the median of the confidences as an alternative score for the whole detection. MediaPipe could not be evaluated as it does not provide scores or confidences. Using its implementation within the Python package scipy.stats, we measured correlations with Spearman Rank Order Coefficients.

### v4

@@ Abstract
Automatic markerless estimation of infant posture and motion from ordinary videos carries great potential for movement studies ”in the wild”, facilitating understanding of motor development and massively increasing the chances of early diagnosis of disorders. There is rapid development of human pose estimation methods in computer vision thanks to advances in deep learning and machine learning. However, these methods are trained on datasets that feature adults in different contexts. This work tests and compares seven popular methods (AlphaPose, DeepLabCut/DeeperCut, Detectron2, HRNet, MediaPipe/BlazePose, OpenPose, and ViTPose) on videos of infants in supine position and in more complex settings. Surprisingly, all methods except DeepLabCut and MediaPipe have competitive performance without additional finetuning, with ViTPose performing best. Next to standard performance metrics (average precision and recall), we introduce errors expressed in the neck-mid-hip (torso length) ratio and additionally study missing and redundant detections, and the reliability of the internal confidence ratings of the different methods, which are relevant for downstream tasks. Among the networks with competitive performance, only AlphaPose could run close to real time (27 fps) on our machine. We provide documented Docker containers or instructions for all the methods we used, our analysis scripts, and the processed data at https://hub.docker.com/u/humanoidsctu and https://osf.io/x465b/.
Introduction
@@
Methods for automatic human keypoint extraction (for example, eyes, wrists, hips, feet, etc.), also called human pose estimation, from images and videos are rapidly evolving, with performances increasing every year (see [34, 35, 36] for recent surveys). These provide key enabling technology to make automatic motion extraction and analysis from “in the wild” recordings possible. These methods were developed primarily to estimate posture and movement of adult bodies (see [37] for a recent comparison of OpenPose and marker-based motion capture). So far, these methods have been applied to infant videos “as is”. However, the morphology of the infant body is different from the body proportions of the adult, especially in early infancy [38, 39]. Additionally, small infants are typically in supine position (on their back) and move differently than standing adults, which constitutes a dataset which is out of the distribution the adult pose estimation models were trained on. Although automated pose estimation seems to be generally good (see Fig. 1 a-c), in a previous study [11], some limitations of current pose estimation methods in infants were highlighted. For example, OpenPose [28] was found to struggle with some camera angles, body postures, and keypoints. In particular, complex leg positions, such as when legs are crossed in supine position, are not often estimated with high accuracy. This can be seen in Fig. 1 (d): OpenPose misses the left ankle keypoint and places the left knee keypoint on the thigh, very close to the left hip. In the current study, comparing different pose estimation methods, we found similar examples. Fig. 1 (e) shows that HRNet Bottom-Up fails on a complex leg position and fails to place the left shoulder and elbow keypoints, placing them on the right; while Fig. 1 (b) shows that HRNet Top-Down hallucinates a second person in the image (see the additional bounding box at the infant’s right forearm), and Fig. 1 (f) shows ViTPose not detecting the infant while sitting on the lap of a parent (the image was cropped to not show the parent’s face in the figure).
@@
Figure 1: Examples of successful and erroneous keypoint estimates obtained from the different human pose estimation methods: a) HRnet Bottom-Up (“Supine” dataset, infant AA, 17 weeks) successful; b) HRnet Top-Down (“Supine” dataset, TH, 10 weeks) with a successful estimate but a second hallucinated person; c) ViTPose (SyRIP image 293) successful; d) OpenPose missing the left ankle keypoint (synthetic infant 1); e) HRNet Bottom-Up failing to estimate a complex leg position (synthetic infant 10); f) ViTPose detecting the parent but not the infant (“Lap” dataset).
@@
Thus, our first goal and contribution is to assess the performance of 2D human pose estimation methods when applied specifically to infants in supine position. In order to provide a more general outlook, we additionally present results on images of infants in a wider variety of postures, including standing or turning away from the camera, as well as with infants on the lap, or sitting or standing in front, of their parent. Only few studies have partially addressed this problem. However, if present, “in the wild” recordings did not constitute the majority of the datasets. Needham et al. [19] used a multi-camera setup in the lab and compared OpenPose, AlphaPose and DeepLabCut. Groos et al. [40] created a large proprietary dataset of human-annotated infant images from both controlled and standardized environments and “in the wild” recordings, and used them to retrain several models, including OpenPose. Their best model achieves on average 81.11% of its estimates being located within an error of 10% of the length of the infant’s face with an average error at 6.8% of the face length, while the “vanilla” OpenPose had 49.66% of its estimates within this margin of error with an average error of 14.3% of the face length. Sermpon et al. [41] have found weak correlations between pose estimation features derived from MediaPipe’s estimates and manually-annotated movements towards the midline. Jahn et al. [42] compared OpenPose, MediaPipe, HRNet, AggPose and VitPose, as well as fine-tuned versions of ViTPose and HRNet (AGMA-HRNet48) on infant recordings in a laboratory context made to specifically target GMA settings (infants in supine position, seen from above and at a diagonal angle). Their results show that ViTPose-huge has the second best results (higher than AGMA-HRNet48, the HRNet model fine-tuned with infant images from a similar GMA context), locating 84.6% of its estimates within 2.5cm of the ground truth (estimated as 10% of the torso length), while their re-trained ViTPose locates 93.89% of estimates within this distance but seems to overfit and to have lower performance than the non-trained ViTPose on other datasets according to their supplementary materials.
@@
In this article, we compare the performance of seven methods: Alphapose [31], DeepLabCut / DeeperCut (v2.2.1.1) [43], Detectron2 (v0.6) [44], MediaPipe / BlazePose (v0.10.14) [45, 46], HRNet (MMPose v0.28.0) [47, 48] Bottom-Up (HRnet BU) and Top-Down (HRnet TD), OpenPose [28], and ViTPose (MMPose v1.1.0) [49].
@@
As a second contribution of this work, we put together and evaluated a number of measures to assess the performance of the pose estimation methods. Besides the ones typically used by the machine learning community (Average Precision and Recall), we consider alternative measures that express the errors scaled to the infant’s body dimensions (Neck-MidHip error), and we additionally analyze the percentage of missing and redundant detections and the reliability of the confidence estimates that the pose estimation methods output. We also report the processing time of the methods. All of these together provide a more complete picture and allow the user to make an optimal choice of the method.
@@
The third contribution of this work constitutes in making all the pose estimation methods and tools for the results analysis available to the community. Although the pose estimation algorithms themselves are publicly accessible in their corresponding repositories, we make available the versions we used including the complete environment in public Docker containers (at https://hub.docker.com/u/humanoidsctu) with instructions and additional abstractions to simplify their use. In addition, we also share the evaluation scripts and the detailed results (at https://osf.io/x465b/).
@@ Methods
To ensure that comparisons are as fair as possible, we used versions of the pose estimation methods that were trained with the same dataset: COCO [50], although some methods only have available weights that used other datasets in their training, instead of or in addition to COCO (see each method’s respective subsection below or the summary Tab. 1). The methods provide 17 estimated keypoints, with the exception of DeepLabCut, MediaPipe and OpenPose that provide 14, 33 and 18 keypoints respectively. We will call a single set of estimated keypoints a detection. The details of the parameters with which the methods were run are provided in the Supplementary Materials. All methods except MediaPipe provide a form of confidence values as their own internal evaluation of the estimated result, with higher values indicating higher confidence in the quality of the estimates. This value can be provided for the whole set of keypoints or for individual keypoints (or both). Following the trend in the documentation and methods’ outputs, we will use score when referring to the whole detection value and confidence when referring to the individual keypoints values. The methods follow either one of the two approaches: Top-Down, which first uses a detector to delimit areas of interest in the image that encompass a person to detect and then estimates a set of keypoints within this space; and Bottom-Up, which finds keypoints or sets of keypoints separately on the whole image before joining them into full keypoint sets and separating them into several individuals if necessary.
@@
  • Real infants “Supine” dataset. 720 annotated images (90 per video) from 8 videos (59200 images total) of two infants, followed longitudinally between 2 and 6 months of age. The annotations were made using the DeepLabCut labeling tool. 480 images (60 per video) were chosen by taking the first ten images every 100 images, from 0 to 9, then 100 to 109, until 509. The remaining 240 images (30 per video) were selected through DeepLabCut’s integrated selection tool using k-means clustering to pick the least similar images in each video. Fourteen keypoints were annotated, corresponding to the COCO keypoints minus the ears and the nose. Our dataset is entirely comprised of infants recorded at home by their parents, each video being different from the others in terms of camera angles, lighting conditions, as well as background and clothes colors and type.
  • Synthetic infants – MINI-RGBD. Based on recordings of real infants, Hesse and colleagues have trained a model capturing infant shape and posture distributions which they subsequently used to produce and render synthetic sequences of artificial infants under 7 months in supine position—the MINI-RGBD dataset [54]. There are 12 synthetic infants with 1000 images each. Ground truth for 25 keypoints is available. Of these 25 keypoints, 13 are in common with COCO’s 17 keypoints and were used to evaluate the methods. Each of the twelve synthetic sequences is unique, including backgrounds with large variations between sequences (see Fig. 1 (d, e) for two examples, or see the original paper [54]).
@@
  • Real infants on the “Lap (sitting)” dataset. 359 annotated images from 6 videos of 6 different infants aged from 17 weeks to 18 months. Two of these videos (infants) were recorded at home by the parents, with the infant sitting on the lap of a parent, themselves sitting on a couch. The remaining four videos (infants) were recorded in a laboratory context, with the camera placed in front of the participants, who were sitting (or standing) on a mat placed on the ground. The parent, and in particular their face, was not always fully visible in the images. The number of 359 images was reached after filtering down the images that were originally annotated but where the infant was missing (moved out of the field of the camera), or because an experimenter was in front of the camera and occluded most of the infant.
@@
For real infants, it is difficult to estimate the difficulty of each sequence or image, though it is clear that our “Lap” and the SyRIP datasets would be more difficult to estimate than our “Supine” dataset. The SyRIP dataset contains a wider variety of infant postures and camera angles, including many that partially hide the infants (e.g., from the side), making the estimation more difficult. As for the “Lap” dataset, since the methods were all originally trained on datasets of adult images, they are biased toward detecting adults, and the presence of the adult overlapping right behind the infant is a strong source of disturbance. For synthetic infants, the MINI-RGBD dataset has a definition of the sequence difficulty: easy (IDs 1 to 4), medium (5 to 9) and difficult (10 to 12), with this difficulty judged primarily with regard to the posture of the infant. Although the main manuscript focuses on the average results, it might be interesting to look at the results at the level of individual videos or their estimated difficulty. Such results are available in the Supplementary Materials.
@@
Our supplementary materials include the results for an extended version of our “Supine” dataset: 720 supplementary images from 8 additional videos; as well as a continuous sequence of 900 images (30 seconds) of a single infant at 17 weeks old. A summary table of each dataset’s characteristics is available in Sup. Tab. 1.
@@
The metrics used for comparison are described below. Because we observed differences in the results when the methods were given input in the form of videos or images, we present the results for both input types independently, except for DeepLabCut and ViTPose (within the MMPose environment) that only accepted video input.
@@
Out of all these metrics, the ones that can be used even without ground truth, and for which we have the results on the full videos on our “Supine” dataset are: percentages of missing data and redundant detections, and the processing speed.
@@
To evaluate the performance of pose estimation, the gold standard metrics in the literature are Average Precision and Average Recall [56], ranging between 0 and 100, higher values indicating better performance. We follow the benchmark evaluation set by the COCO challenge to calculate AP and AR, averaging the values computed over ten thresholds from 0.5 to 0.95, with steps of 0.05. Roughly, if the OKS for a detection is above the threshold and there is any ground truth that has not been matched with a detection yet on a given image, then the detection counts as True Positive. Other cases count as False Positives. If there is no detection but there is a ground truth, then it counts as a False Negative. Precision and Recall are computed for each threshold, then averaged. Note that despite the name Average Precision, it is actually computed as the area under the Precision-Recall Curve [57, 56]. In our case, the computation is simplified, as all our annotated images contain one single infant (one ground truth). For our “Lap” dataset, the ground truth annotations of the parents were ignored, so that each method was only evaluated against the infant annotations.
@@
As an alternative to OKS, we additionally use a reference distance, the Neck-MidHip length, corresponding roughly to the length of the infants’ torso. With this reference distance, we normalize the errors between recordings to compensate for different camera settings and infants of different ages and with different body sizes. Unlike OKS, such errors are easier to grasp as they directly relate to a measurable part of the body of the infant. Lower is better, and the minimum is 0. Others and ourselves have previously used a similar strategy in a similar context [42, 11, 29, 27].
The head length used in [40] can be an alternative. However, the torso is a more stable body part that does not rotate as much as the head, especially for infants in supine position, making it a better choice as a normalization reference. Additionally, neither the ground truth following the COCO format [50] nor the MINI-RGBD ground truth [54] contain keypoints that can be used to compute the length of the head without larger approximations than for the torso.
The equations are shown in Eq. 2. Given $j$ a recording with $I_{j}$ annotated images for which Neck and MidHip keypoints ground truth are available or can be approximated. Given the images from the recording $i\in[0,I_{j}]$, then the Neck-MidHip length for this image $nmh\_len_{ji}$ is defined in pixels as the Euclidean distance between the Neck and the MidHip ground-truth coordinates in this image. For our “Supine” and the MINI-RGBD dataset, we take the median of all $nmh\_len_{ji}$ as the NeckMidHip length to use to normalize the recording: $nmh\_len_{j}$. Given a keypoint $k$ that is visible in the image and annotated, and detected by the pose estimation method, then its Neck-MidHip error $nmh\_error_{kij}$ is its error (Euclidean distance) relative to its ground truth, $d_{kij}$, divided by the median Neck-MidHip length for the recording, $nmh\_len_{j}$. For our “Lap” and the SyRIP dataset, which contain images from different sources mixed together, we normalize using each images’ neck-MidHip distance $nmh\_len_{ji}$ instead of a median from a whole recording. This has the downside that only images where this length can be estimated are used to calculate the error. Then, we obtain per-keypoint-location mean Neck-MidHip errors, and an overall mean full-body error, from all images and recordings.
@@
For the synthetic infants in the MINI-RGBD dataset, the Neck and MidHip positions are provided directly in the ground truth. In the real infants datasets, the ground truths do not provide Neck or MidHip keypoints. Instead, the centre of the right and left Shoulder and Hip keypoints are used, turning into a “MidShoulder-MidHip” length. We will simply write Neck-MidHip for simplification in the rest of the paper. On our “Supine” dataset and on the MINI-RGBD dataset, the Neck-MidHip length is computed using the median of the Neck-MidHip lengths for a whole video. This enables the use of all annotated images and keypoints even if one of the keypoints needed to compute the length is missing from the ground truth in a particular image (occlusions). It is also less subject to noise in the annotations, with the drawback that a single length is used for a whole video and does not fit exactly to each individual image. Using the median can only be done when there are no changes in camera settings, in particular zoom or camera position, which can happen, for example, when recordings are performed at home by a parent holding a smartphone with their hand; otherwise, it requires to separate the recording into subsets according to the different settings, or use per-image normalization instead, which is what we opted to do for our “Lap” and the SyRIP datasets.
@@
The total percentage of missing data computation is described in Eq. 3. We define the maximum amount of data $maxdata_{m}$ that can be estimated by a method $m$ as the number of images to estimate $N$ times the number of keypoints provided by the method $m_{kp}$. Missing detections mdetmdet are converted to this unit by multiplying with $m_{kp}$ as well, since any missing detection is equivalent to missing all keypoints in this one image. This number is added to the count of individual missing keypoints mkpmkp, and divided by $maxdata_{m}$ to obtain the total percentage of missing data for this method: $mdata_{m}$.
@@
Some methods can erroneously detect more people than there are in the image, as shown in Fig. 1 (b), where an additional ”person” is detected. These redundant detections might indicate that the method hallucinates other people on the image where there are none, or provides several estimates for the same person with different positions. Such behavior is undesirable, in particular for applications with more complex environments that, for example, involve interaction with an experimenter or the parents, such as in our “Lap” dataset, as it complexifies tracking each separate individual when ground truth is not available.
The percentage of redundant detections is computed as the ratio between i) the difference in the number of detections between the expected and provided detection numbers (number of estimated detections minus the number of expected detections) divided by ii) the number of images the method provided detections for (total number of images minus the number of missing detections). For example, if there are 150 detections for 100 input images with 100 expected detections (1 infant per image), then 50 detections would be considered redundant, resulting in a percentage of 50% of redundant detections, translating that, on average, the method provides one additional detection every two images where it detected something. A percentage of 100% would mean that, on average, the method always provides two detections for each image, when one is expected. DeepLabCut, MediaPipe, and OpenPose did not have redundant detections (MediaPipe and OpenPose, have a built-in parameter to limit the detections to 1 person per image).
In the case of the “Lap” dataset, where an adult is always present behind the infant, there are also cases where only the adult and not the infant is detected (Fig. 1 (f)). We empirically observed that some pose estimation methods have a tendency to often detect only the adult and not the infant in such context. This would show as a negative ratio of redundant detections. Taking the example from above, if a method had provided 150 detections for 100 images where 200 detections were expected due to the presence of both an infant and an adult, then we would get -50% “redundant” detections: on average, that method provided the two expected detections for half of the images, and only provided one detection for the other half. For this particular dataset, this is not a foolproof measure, as it is possible that a method detected the adult and then provided a second, redundant, detection of the adult or hallucinated another person somewhere else in the image instead of detecting the infant.
@@
In a real application scenario, without available ground truth, confidence or score values can be useful to provide cues about detections or keypoints that should not be trusted. For example, confidences are used as coefficients in the optimization process of smplify-x [58] when optimizing the parameters of the 3D full-body pose estimation from the 2D pose estimation. These values should be positively correlated with the OKS (the higher the OKS, the higher the score). When available, we use the score provided by the methods for each image. When not available, we used the median of the confidences as an alternative score for the whole detection. MediaPipe could not be evaluated as it does not provide scores or confidences. Using its implementation within the Python package scipy.stats [59], we measured correlations with Spearman Rank Order Coefficients.